ln (x+5) - ln (x-1) = ln (3)
Answer by stanbon(48569)

ln (x+5) - ln (x-1) = ln (3)
ln[(x+5)/(x-1)] = ln3
Take the anti-ln of both sides to get:
[(x+5)/(x-1)] = 3
Cross-multiply to get:
x+5 = 3x-3
2x=8
x=4
============
Cheers,
Stan H.